How do you convert 7.5ft.squared to inches squared

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

How do you convert 7.5ft.squared to inches squared
=============
I convert it very well, thx for asking.
----
1 ft = 12 inches
1 ft*1 ft = 12 inches * 12 inches.
(1')^2 = (12")^2
=======================================
1' = 12"
(1')^3 = (12")^3 (in case you wanna see how that works)

Answer by ikleyn(52781)

.

HINT

        (1)   1 foot = 12 inches.

        (2)   1 square foot = 12*12 square inches = 144 square inches.

        (3)   7.5 square feet is 7.5 times one square foot.

        (4)   It is your turn.